# Pitfalls and Outlooks in Using COMET

### v2

@@ Abstract
Since its introduction, the COMET metric has blazed a trail in the machine translation community, given its strong correlation with human judgements of translation quality. Its success stems from being a modified pre-trained multilingual model finetuned for quality assessment. However, it being a machine learning model also gives rise to a new set of pitfalls that may not be widely known. We investigate these unexpected behaviours from three aspects: 1) technical: obsolete software versions and compute precision; 2) data: empty content, language mismatch, and translationese at test time as well as distribution and domain biases in training; 3) usage and reporting: multi-reference support and model referencing in the literature. All of these problems imply that COMET scores is not comparable between papers or even technical setups and we put forward our perspective on fixing each issue. Furthermore, we release the SacreCOMET package that can generate a signature for the software and model configuration as well as an appropriate citation. The goal of this work is to help the community make more sound use of the COMET metric.
@@ 1 Introduction
Automated metrics provide a cheap and scalable way of evaluating and benchmarking NLP models. In machine translation (MT), the evaluation protocol has moved from string matching metrics (BLEU, TER, chrF, inter alia; Papineni et al., 2002; Snover et al., 2006; Popović, 2015) to trained metrics based on neural networks (Shimanaka et al., 2018; Takahashi et al., 2020; Rei et al., 2020a; Sellam et al., 2020) with COMET being the most widely used option in the community. The trained metrics have been shown to correlate much better with human judgement (Freitag et al., 2021, 2022b, 2023), making them more reliable in assessing or estimating translation quality and ranking translation systems.
Nonetheless, the solution to translation evaluation is yet to be perfected. One problem is the misuse and misunderstanding of the metric. Previously, Post (2018) showed that different usages and implementations of BLEU, e.g. tokenization and smoothing, can lead to vast inconsistencies in reported scores. We suspect that sensitivities can arise from several stages of using neural metrics too—e.g. obtaining COMET involves running a particular checkpoint with the COMET software on some particular hardware. This process might be sensitive to misconfigurations in each stage, resulting in unexpected or nonreproducible behaviours. Furthermore, trained MT metrics like COMET often fit on limited data intended to be valid translations, leading to reduced robustness owing to its nature as an optimized machine learning model.
In this work, we
list nine problems covering technical issues, data bias, and model reporting;
we show that inconsistent use of COMET leads to scores non-comparable across papers or setups;
release SacreCOMET package for better reporting and reproducibility.
@@ 2 Background and Setup
Human judgements for translation quality come from annotation campaigns, where translations are evaluated with some annotation protocol. Historically for WMT, the protocol was a variant of direct assessment (DA; Graham et al., 2013) which has annotators providing a number from 0 (lowest) to 100 (highest) as the overall segment quality. This has been subsequently replaced by MQM and ESA protocols (Lommel et al., 2014; Kocmi et al., 2024c), though DA remains the most abundant data source for neural metric training.
Automated metrics aim to yield scores that correlate with human judgements of translations. Most metric scores are computed at the segment level and then aggregated at the system level. The evaluation of metrics is done with respect to the human judgements.
@@
Metrics such as chrF or BLEU are heuristic algorithms that match n-grams between the translation and the reference to compute a score. In contrast, COMET is a machine learning model fine-tuned from a pre-trained multilingual language model, e.g. XLM-R (Conneau et al., 2020), with an additional regression head. A reference-based COMET model learns to regress from a tuple of [source, hypothesis, reference] to the human judgement score (from previous evaluation campaigns) at the segment level. The quality estimation version of COMET is prepared by omitting the reference in the input.
Most issues in this work are demonstrated using two COMET checkpoints unless noted otherwise: the reference-based COMET and the reference-free COMET (Rei et al., 2022a).111huggingface.co/Unbabel/wmt22-comet-da,222huggingface.co/Unbabel/wmt22-cometkiwi-da Both metrics output to a normalized range between 0 and 1. The COMET framework unbabel-comet is of version 2.2.2 except when we test how different software versions affect COMET scores.
@@ 3 Problems
Tool usage and score interpretation: it has no defined way of equipping multiple references when available which leaves room for research. From a bibliometric perspective, we reveal that some literature omits a clear reference to the checkpoint version or citation.
@@ 3.1 Software versions [technical]
The official installation of the COMET package requires Python 3.8 or above.333github.com/Unbabel/COMET 332dfb0 as of Aug 2024. We demonstrate that neglecting this would lead to unexpected scores because the same COMET checkpoint can produce vastly different scores with previous COMET framework versions that are no longer supported.
Under several Python versions, executing the following code leads to different COMET package (unbabel-comet) versions being installed. Running the framework for translation evaluation will subsequently result in false conclusions as shown in Table 1’s evaluation on WMT23 tests. The direct cause is that the discontinued Python 3.7 supports only up to unbabel-comet 1.1.2. Nonetheless, we caution that the underlying factor is the version of unbabel-comet rather than Python.
@@
Updating both Python and unbabel-comet to the latest version is helpful and reporting the version of the toolkit can aid reproducibility. The package version should also be reported.
| Precision | COMET | MAE | Acc | sec | ||
@@ 3.2 Numerical precision [technical]
Despite the aforementioned advantages, model quantization is not a feature supported by the current COMET framework by default. We thus make minimal modifications to the software and investigate the effect of numerical precision on COMET scores on both CPU (FP32, FP16, and QINT8) and GPU (FP32 and FP16). For CPU inference with FP16 or dynamic QINT8, we apply the quantization module torch.ao.quantization from PyTorch before model.predict(). When using GPU-FP16, we first load the model weights to FP32. After that, we call .half() to explicitly convert the weights of the encoder, i.e., XLMRobertaModel() from transformers. This is because we discovered that loading the weights directly in FP16 would still result in FP32 precision.
@@
| EnDe | EnZh | DeEn | ZhEn | EnDe | EnZh | DeEn | ZhEn | |
| Baseline (Online-A) | 0.837 | 0.862 | 0.855 | 0.803 | 0.800 | 0.791 | 0.794 | 0.787 |
| All empty | 0.335 | 0.392 | 0.353 | 0.374 | 0.315 | 0.319 | 0.537 | 0.467 |
| Random sent. | 0.373 | 0.434 | 0.334 | 0.350 | 0.333 | 0.341 | 0.447 | 0.391 |
| Random sent. w/ words shuffled | 0.244 | 0.419 | 0.264 | 0.347 | 0.232 | 0.325 | 0.307 | 0.385 |
We use NVIDIA GeForce RTX 3090 for GPU inference and a batch size of 8 in all settings (in practice a quantized model makes room for a larger batch size). Table 2 summarises the effect of numerical precision. In addition to reporting COMET scores, we also report 1) inference time in seconds (sec) as an efficiency measure; and 2) segment-level mean absolute error (MAE), segment-level Kendall’s tau-c (), and system-level pairwise accuracy (Acc). Everything is compared to the human DA scores either on segment- or system-level. Technically, Kendall’s calculates rank correlation on an ordinal scale with adjustments for ties and pairwise accuracy computes the proportion of system pairs that have the same ordering by a metric as by humans.
Our results show that there is no meaningful difference between FP32 and FP16 in both CPU and GPU devices up to 3 significant figures. On GPU, FP16 is about 30% faster in time, but unsurprisingly it does not provide any speed-up on CPU. Interestingly, on the CPU, dynamic QINT8 gives systematically higher COMET scores and shorter running time than FP32 and FP16. However, the much lower and pairwise accuracy indicate the lack of reliability at this precision. In addition to precision, we explored the effect of batch size and the choice of GPU or CPU during inference with results listed Appendix A. Whilst there are some fluctuations, they are mostly negligible. However, lower precision allows for higher batch size which usually directly corresponds to speed-up.
@@ 3.3 Empty hypothesis [data]
An empty translation (string of length 0) gets penalized heavily by string-based metrics because an empty string has zero surface overlap with the reference. However, neural metrics provide no such guarantee. We show that COMET assigns a positive instance-level score even if the hypothesis is an empty string as corroborated by Lo et al. (2023). Problematically, this score can even occasionally be higher than the score for a genuine system translation.
In Table 3, we list COMET scores for system Online-A’s hypotheses at WMT23 and a file full of empty lines. Furthermore, we compare them with completely incorrect translations to explicate the score magnitude in two ways. 1) Random sent.: we shuffle at the sentence level WMT22’s reference files in the respective translation directions. This provides us with high-quality human-written sentences. We sub-sample or over-sample if the number of lines in WMT22 is larger or smaller than the WMT23 size. 2) Random sent. w/ words shuffled: we further shuffle the words at each line in the sentence-shuffled files, generating nonsensical sentences. Sentence-shuffled hypotheses can be seen as fluent but extremely inadequate sentences whereas word-shuffled sentences are neither fluent nor adequate.
@@
We then count the number of empty lines that score better than a translation from Online-A in Table 4. We observe roughly 0.25% such cases for most translation directions except for DeEn’s COMET score situating at 1.45%. Further, in Figure 1 we plot the distributions of COMET scores for ZhEn’s empty and genuine translations with other translation directions in Appendix B. Noticeable overlaps are observed for COMET when translating into English.
Adhering to the guidelines of the DA protocol, this is not the correct behaviour because an empty hypothesis should receive a score of 0, to match the category no meaning preserved. This, however, is not surprising with COMET, because it has likely not seen empty hypotheses during training that would have received a score of 0 from a human annotator. Finally, since the distributions of empty hypotheses are nowhere close to a single vertical bar, it implies that segment-level COMET scores oddly hinge on the source sentence, as noted by Sun et al. (2020).
@@ 3.4 Hypothesis language mismatch [data]
String overlap-based metrics can also score a hypothesis in a language different from the reference almost zero. However, even for the reference-based COMET, there is no explicit way to enforce the intended target language given it is an embedding-based metric. In the field of translation and translation evaluation, this poses an increasingly pronounced problem, especially for multilingual translation models as well as the recent large language models, in which the generated language cannot be as easily controlled as it could be in encoder-decoder models (Zhang et al., 2023).
We conduct experiments to understand if translation outputs in an incorrect language leave an impact on the score, and whether different mismatching languages can lead to distinct patterns. We use the translation directions EnRu, EnUk, and EnZh in WMT23 which share the same English source input. Having Online-A’s output in all three directions, we substitute hypotheses in a particular translation direction with those from another direction. A similar hypothesis was presented in Amrhein et al. (2022) which suggested that COMET metrics are not robust to hypothesis language mismatch. Our experiment setup offers a more detailed evaluation setup than their contrastive setup.
Table 5 presents COMET scores for translations in correct and incorrect languages, as well as empty lines and random sentences in the correct language as “baselines”, deemed as completely wrong translations. The pattern shows that when the hypothesis is in a language distant from the reference, COMET score declines much more than when the hypothesis is in a more similar language. A more concerning issue is that even when the hypotheses’ language is completely wrong, the resulting COMET score can still be vastly higher than empty hypotheses or random sentences in the correct language. We omit COMET because it does not have a mechanism to read a reference (language) making it inherently incapable of distinguishing output languages.
@@
Run language identification and set hypotheses in an unexpected language to have 0 COMET score before aggregating them system-level. Also, check with a string overlap-based metric like BLEU or chrF.
| Mismatch () Hyp. lang () | EnRu | EnUk | EnZh | |||
| Uk | Zh | Ru | Zh | Ru | Uk | |
| 100% | 0.797 | 0.536 | 0.807 | 0.540 | 0.655 | 0.644 |
| 0% | 0.853 | 0.832 | 0.862 | |||
| All empty | 0.316 | 0.329 | 0.391 | |||
| Random sent. | 0.463 | 0.472 | 0.435 | |||
3.5 Score distribution bias [data]
@@
We first verify this in Table 6 which shows that WMT translation directions are associated with vastly different human DA scores (from 0.51 to 0.91). Empirically, we illustrate this issue using two high-resource directions EnDe and EnZh. As shown in Figure 2, we keep either the top 75% or bottom 75% of all scored translations to alter the score distribution for each translation direction, simulating the scenario where low- and high-performing system submissions are received for different directions. We then train different COMET models on the data before and after alteration as per Figure 2. Finally, we evaluate those checkpoints on the same test set and report results in Table 7.
As expected, for both EnDe and EnZh, training on the top or bottom-scoring data would lead to increased or decreased COMET scores on the same set of hypotheses. In addition, we observe that altering scores in a particular translation direction would not affect scores in another direction. For example, removing the bottom 25% EnDe scores “improves” the test-time score from 0.770 to 0.790 whilst EnZh remains unaffected.

### v3

@@ Abstract
The COMET metric has blazed a trail in the machine translation community given its strong correlation with human judgements of translation quality. Its success stems from being a pre-trained multilingual model finetuned for quality assessment. However, it being a neural metric also gives rise to a set of pitfalls that may not be widely known. We investigate these unexpected behaviours from three aspects: 1) technical: obsolete software versions and compute precision; 2) data: empty content, language mismatch, and translationese at test time as well as distribution and domain biases in training; 3) usage and reporting: multi-reference support and model referencing in the literature. All of these problems imply that COMET scores may be not incomparable between papers or technical setups and we put forward our perspective on fixing each issue. Furthermore, we release the sacreCOMET package that can generate a signature for the software and model configuration as well as an appropriate citation. The goal of this work is to help the community make more sound use of the COMET metric.
@@ 1 Introduction
Automated metrics provide a cheap and scalable way of evaluating and benchmarking NLP models. In machine translation (MT), the evaluation protocol has moved from string matching metrics (BLEU, TER, chrF, inter alia; Papineni et al., 2002; Snover et al., 2006; Popović, 2015) to trained neural metrics (Shimanaka et al., 2018; Takahashi et al., 2020; Rei et al., 2020a; Sellam et al., 2020) with COMET being widely adopted. The trained metrics have been shown to correlate much better with human judgement (Freitag et al., 2021, 2022b, 2023), making them more reliable in estimating translation quality and ranking translation systems.
Nonetheless, the solution to translation evaluation is yet to be perfected. One problem is the haphazard use of the metric. Previously, Post (2018) showed that different usages and implementations of BLEU, e.g. tokenization and smoothing, lead to inconsistencies in scores. We suspect that the use of COMET might be sensitive to misconfigurations too, resulting in unexpected behaviours. Furthermore, trained MT metrics are optimized on a limited amount of data (usually valid machine translations), leading to overfitting and reduced robustness against corner cases. Contributions of this work are listed as follows:
we reveal nine problems spanning technical issues, data biases, and model reporting;
we show that inconsistent use of COMET leads to non-comparable scores across papers or setups;
we release the sacreCOMET package for better reporting and reproducibility;
we provide directions for future work on building learned metrics.
@@ 2 Background and Setup
Publicly available human judgements of translation quality come from shared task annotation campaigns, where translations are evaluated with some annotation protocol. From 2017, in WMT, the protocol was a variant of direct assessment (DA; Graham et al., 2013) which has annotators providing a number from 0 (lowest) to 100 (highest) as the segment quality. This has been subsequently replaced by MQM and ESA protocols (Lommel et al., 2014; Kocmi et al., 2024d), though DA remains the most abundant data source for neural metric training.
Automated metrics aim to yield scores that correlate with human judgements of translations. Most metric scores are computed at the segment level and then aggregated at the system level to e.g. obtain system comparison. The evaluation of metrics is done with respect to the human judgements.
@@
Metrics such as chrF or BLEU are heuristic algorithms that match n-grams between the translation and the reference to compute a score. In contrast, COMET is a machine learning model fine-tuned from a pre-trained multilingual language model, e.g. XLM-R (Conneau et al., 2020), with an additional regression head. A reference-based COMET model learns to regress from a tuple of [source, hypothesis, reference] to the human judgement score (from previous evaluation campaigns) at the segment level. The quality estimation (reference-free) version of COMET is prepared by omitting the reference from the input.
Most issues in this work are demonstrated using two COMET checkpoints unless noted otherwise: the reference-based COMET and the reference-free COMET (Rei et al., 2022a). Both metrics output to a normalized range between 0 and 1. The COMET framework unbabel-comet is of version 2.2.2 except when we test how different software versions affect COMET scores.
@@
We base our experiments on the general domain translation and metrics shared tasks of WMT from 2023 (Kocmi et al., 2023; Freitag et al., 2023). The translation directions in the paper are centred around EnDe and EnZh, though we occasionally include other translation directions for demonstrative purposes.
Whenever possible, we divide all scores, including DA and model outputs, such that their output is between 0 and 1.
@@ 3 Problems
Tool usage and score interpretation: COMET has no defined way of equipping multiple references when available which leaves room for research. From a bibliometric perspective, we reveal that some literature omits a clear reference to the checkpoint version or citation.
@@ 3.1 Software versions [technical]
The official installation of the COMET package requires Python 3.8 or above.111github.com/Unbabel/COMET 332dfb0 as of Aug 2024. We demonstrate that neglecting this would lead to unexpected scores because the same COMET checkpoint can produce vastly different scores with previous COMET framework versions that are no longer supported.
Under several Python versions, executing the following code leads to different COMET package (unbabel-comet) versions being installed. Running the framework for translation evaluation will subsequently result in false conclusions as shown in Table 1’s evaluation on WMT23 tests. The direct cause is that Python 3.7, which has been discontinued, only supports unbabel-comet versions up to 1.1.2. Nonetheless, we caution that the underlying factor is the version of unbabel-comet rather than Python.
@@
Updating both Python and unbabel-comet to their latest versions is helpful and reporting the toolkit version aids reproducibility.
| Precision | COMET | MAE | Acc | Time (s) | ||
@@ 3.2 Numerical precision [technical]
Despite the aforementioned advantages, model quantization is not a feature supported by the current COMET framework by default. We thus make minimal modifications to the software and investigate the effect of numerical precision on COMET scores on both CPU (FP32, FP16, and QINT8) and GPU (FP32 and FP16). When using FP16, we first load the model weights to FP32, followed by .half() call. This is because loading the weights directly in FP16 still incorrectly results in FP32 precision. For CPU inference with dynamic QINT8, we apply the quantization module torch.ao.quantization from PyTorch.
@@
| Hypothesis () | EnDe | EnZh | DeEn | ZhEn | EnDe | EnZh | DeEn | ZhEn |
| Real system (Online-A) | 0.837 | 0.862 | 0.855 | 0.803 | 0.800 | 0.791 | 0.794 | 0.787 |
| Empty hypothesis | 0.335 | 0.392 | 0.353 | 0.374 | 0.315 | 0.319 | 0.537 | 0.467 |
| Random hypothesis (fluent) | 0.373 | 0.434 | 0.334 | 0.350 | 0.333 | 0.341 | 0.447 | 0.391 |
| Random hypothesis (shuffled words) | 0.244 | 0.419 | 0.264 | 0.347 | 0.232 | 0.325 | 0.307 | 0.385 |
We use AMD Ryzen 9 5900X with NVIDIA GeForce RTX 3090 for GPU inference and a batch size of 8 in all settings (in practice a quantized model makes room for a larger batch size). Table 2 summarises the effect of numerical precision. In addition to reporting COMET scores, we also report (1) inference time in seconds (sec) as an efficiency measure; and (2) segment-level mean absolute error (MAE), segment-level Kendall’s tau-c (), and system-level pairwise accuracy (Acc). Everything is compared to the human DA scores either on segment- or system-level. Technically, Kendall’s calculates rank correlation on an ordinal scale with adjustments for ties and pairwise accuracy computes the proportion of system pairs that have the same ordering by a metric as by humans.
Our results show that there is no meaningful difference between FP32 and FP16 in both CPU and GPU devices up to 3 significant figures. On GPU, FP16 is about 30% faster in time, but unsurprisingly it does not provide any speed-up on CPU. Interestingly, on the CPU, dynamic QINT8 gives systematically higher COMET scores and shorter running times than FP32 and FP16. However, the much lower and pairwise accuracy indicate the lack of reliability at this precision. In addition to precision, we explored the effect of batch size and the choice of GPU or CPU during inference with results listed Appendix A. Whilst there are some fluctuations, they are mostly negligible. However, lower precision allows for higher batch size which usually directly corresponds to speed-up.
@@ 3.3 Empty hypothesis [data]
An empty translation (a string of length 0) gets penalized heavily by string-based metrics because an empty string has zero surface overlap with the reference. However, neural metrics provide no such guarantee. We show that COMET assigns a positive instance-level score even if the hypothesis is an empty string as corroborated by Lo et al. (2023). Problematically, this score can even occasionally be higher than that of a genuine system translation.
In Table 3, we list COMET scores for system Online-A’s hypotheses at WMT23 and a file full of empty lines. Furthermore, we compare them with completely incorrect translations to explicate the score magnitude in two ways:
Random hypothesis: we shuffle WMT22’s reference files at the sentence level in the respective translation directions. This provides us with high-quality human-written sentences. We sub-sample or over-sample if the number of lines in WMT22 is larger or smaller than the WMT23 size.
Random hypothesis (shuffled words): we further shuffle the words at each line in the sentence-shuffled files, generating nonsensical sentences.
Sentence-shuffled hypotheses can be seen as fluent but extremely inadequate sentences whereas word-shuffled sentences are neither fluent nor adequate.
@@
We count the number of empty lines that score better than a translation from Online-A in Table 4. We observe roughly 0.25% such cases for most translation directions except for DeEn’s COMET score situating at 1.45%. Further, in Figure 1 we plot the distributions of COMET scores for ZhEn’s empty and genuine translations with other translation directions in Appendix B. Noticeable overlaps are observed for COMET when translating into English.
Adhering to the DA protocol guidelines, this is not the proper behaviour because an empty hypothesis should receive a score of 0, to match no meaning preserved. This however is unsurprising with COMET, which has likely not seen empty hypotheses during training that would have received a score of 0 from a human annotator. Finally, even by relaxing the 0-score expectation, the metric should still assign the same score to all empty hypotheses regardless of the source. Since the distributions of empty hypotheses are nowhere close to a single vertical bar in Figure 1, it exposes the issue that segment-level COMET scores oddly hinge on the source sentence, as noted by Sun et al. (2020).
@@ 3.4 Hypothesis language mismatch [data]
String overlap-based metrics can also score a hypothesis in a language different from the reference almost zero, especially with script mismatch. However, even for the reference-based COMET, there is no explicit way to enforce the intended target language. This poses an increasingly pronounced problem, especially for multilingual translation models as well as the recent large language models, in which the generated language cannot be as easily controlled (Zhang et al., 2023).
We conduct experiments to understand if translation outputs in an incorrect language impact the score, and whether different mismatching languages can lead to distinct patterns. We use the translation directions EnRu, EnUk, and EnZh in WMT23 which share the same English source input. Having Online-A’s output in all three directions, we substitute hypotheses in a particular translation direction with those from another direction. A similar hypothesis was presented by Amrhein et al. (2022) which suggested that COMET metrics are not robust to hypothesis language mismatch. Our experiment setup offers a more detailed evaluation setup than their contrastive setup.
Table 5 presents COMET scores for translations in correct and incorrect languages, as well as empty lines and random sentences in the correct language as “baselines”, deemed as completely wrong translations. The pattern shows that when the hypothesis is in a language distant from the reference, the COMET score declines much more than when the hypothesis is in a similar language. A more concerning issue is that even when the hypotheses’ language is completely wrong, the resulting COMET score can still be vastly higher than empty hypotheses or random sentences in the correct language. We omit COMET because it does not have a mechanism to read a reference (language) making it inherently incapable of distinguishing output languages.
@@
Run language identification and set hypotheses in an unexpected language to have a 0 COMET score before aggregating them system-level. Also, check with a string overlap-based metric like BLEU or chrF.
| EnRu | EnUk | EnZh | |
| Correct lang. | 0.853 | 0.832 | 0.862 |
| Incorrect target lang. | Uk: 0.797 | Ru: 0.807 | Ru: 0.655 |
| Zh: 0.536 | Zh: 0.540 | Uk: 0.644 | |
| Empty hyp. | 0.316 | 0.329 | 0.391 |
| Random hyp. | 0.463 | 0.472 | 0.435 |
@@
3.5 Score distribution bias [data]
We first verify this in Table 6 which shows that WMT translation directions are associated with vastly different human DA scores (from 0.51 to 0.91). Empirically, we illustrate this issue using two high-resource directions EnDe and EnZh. As shown in Figure 2, we keep either the top- or bottom-75% of all scored translations to alter the score distribution for each direction, simulating the scenario where low- and high-performing system submissions are received for different directions. We then train different COMET models on the human train data before and after alteration as per Figure 2. Finally, we evaluate those checkpoints on the same test set and report results in Table 7.
As expected, for both EnDe and EnZh, training on the top or bottom-scoring data leads to increased or decreased COMET scores on the same set of hypotheses. Besides, we observe that altering scores in a particular translation direction inconsistently affects scores in another direction systematically. For example, removing the bottom 25% EnDe scores “improves” the test-time score from 0.770 to 0.790 whilst EnZh remains unaffected.
Finally, an empty translation should have the same score irrespective of the original source. Due to the statistical learning nature of the metric, this is not the case, as found in Section 3.3. Sun et al. (2020) and Zouhar et al. (2023) show that some meaningful correlation with human scores can be attained with just the source as the input. This shows, that there is a learned bias based on the prior difficulty of the source segment, which is undesirable for an objective evaluation metric.